OpenStudy (anonymous):

Which ratio is equivalent to 4 : 28? A. 1 : 8 B. 1 : 7 C. 8 : 52 D. 3 : 14

OpenStudy (anonymous):

4:28 = 4/28 =(4÷4)/(28 ÷ 4) =1/7 =1:7(A)